# Hermitian and non-Hermitian topology in active matter

### v1

@@ 3.1 Band theory and its classical counterpart
[…]
@@
Techniques of the Bloch Hamiltonian and the band structures are also useful in classical physics including active fluid. Basically, the Bloch Hamiltonian corresponds to the matrix obtained via the linear stability analysis of the Fourier modes in fluids [130]. Then, if each Fourier mode is marginally stable, its dispersion relation can be considered as a band structure of fluid. Recent studies [24, 25, 131, 132, 133] have utilized the notions in condensed matter physics to develop metamaterials that control the dynamics of photonic and phononic systems. Metamaterials typically possess a periodic structure of waveguides, which has a clear analogy to the periodic alignment of atoms in crystalline solids.
@@ 3.2.1 Quantum Hall effect
Topology is a property of matter that is invariant under a continuous deformation. For example, considering a genus, the number of holes, we cannot either increase or decrease the genus without cutting the geometry or sticking different surfaces together. Thus, the genus is a topological invariant, and we cannot continuously transform objects with different genera into each other, as exemplified by a ball and a donut. Condensed matter physics employs the notion of topology into a wavenumber space by considering the topology of band structures [15, 16]. To discuss the topology of band structures, one must define the prohibited deformation of the Hamiltonian. In topological insulators, we exclude the deformation under which band structures cross the Fermi energy $E=E_{f}$. Under such continuous deformation (i.e., adiabatic deformation), the bulk is always insulating. Then, one can find defective structures in a wavenumber space, and the number of such defects must be preserved under the continuous deformation in a similar manner as in the preservation of the genus. Thus, insulating phases of matter can be classified by such a topological number, which leads to the notion of topological insulators. Below we shall delineate the definition and physical meanings of the defective structures in band structures.
Figure 4: Quantum Hall effect. (a) Schematic of the quantum Hall effect. Cycrotolon motions of electrons are prevented at the edge of the sample, and instead, the skipping orbit can emerge, which induces the chiral edge current. (b) Experimental results of the quantized Hall voltage. Panel (b) is adapted from K. v. Klitzing, G. Dorda, and M. Pepper, “New Method for High-Accuracy Determination of the Fine-Structure Constant Based on Quantized Hall Resistance” Phys. Rev. Lett. 45, 494 (1980) [17], Copyright ©  1980 by the American Physical Society.
A prototypical example of topological materials is a quantum Hall system, which was experimentally discovered by von Klitzing et al. in 1980 [17]. They measured the Hall conductivity in semiconductors under the existence of the external magnetic field. They tuned the Fermi energy by changing the gate vorticity and obtained the contributions of different Landau levels. If the Fermi energy exists between Landau levels, the longitudinal conductivity becomes zero, and the Hall conductivity takes a constant value until the Fermi energy is increased to the higher Landau level. Thus, the Hall conductivity exhibits plateaus as shown in Fig. 4. Furthermore, the Hall conductivity is proportional to the number of occupied Landau levels and thus is quantized. Thouless et al. [18] revealed that the quantization of the Hall conductance originates from the nontrivial topology of the wave functions. Their theoretical analysis starts from the Hamiltonian of an electron under the external magnetic field,
@@
[…]
Here, we assume that the Fermi energy exists in the band gap. By considering the uniform electronic field $\vec{E}=(E,0)$ as a perturbation and using the Kubo formula, one can calculate the Hall current $\langle j_{y}\rangle=\langle e[\hat{y},H]\rangle$ at zero temperature [134]. This leads to the following expression of the Hall conductance
@@
While the Chern number and the quantized Hall conductance are the bulk properties of topological materials, nontrivial band topology also accompanies unconventional localized modes at the edge of the sample. The appearance of such edge modes can be intuitively understood as follows; since electrons under a uniform magnetic field exhibit cyclotron motions, each electron circularly moves in the bulk of the quantum Hall system. Meanwhile, near the boundary of the sample, the circular motions are interfered by the potential barrier, and instead, electrons bounce at the boundary. The iteration of the curved motions and the bounds leads to a skipping orbit of an electron, where electrons unidirectionally flow along the edge of the sample. Therefore, the quantum Hall system exhibits edge-localized modes that exhibit unidirectional edge flow.
@@
where $\langle\mathbf{r},\mathbf{r}^{\prime}\rangle$ and $\langle\langle\mathbf{r},\mathbf{r}^{\prime\prime}\rangle\rangle$ represent the pairs of nearest and next-nearest neighbor sites, respectively, and ${\rm H.c.}$ represents the Hermitian conjugate of the previous term. $t_{1}$, $t_{2}$ are real-valued parameters determining the hopping amplitudes, and $\phi$ is the phase gained via the hopping to the next-nearest site. $\sigma(\mathbf{r})$ determines the sign of the on-site potential and depends on the sublattice. Such phase factor plays the role of a vector potential (in other words, a gauge field) in a lattice system, and the direction of the effective vector potential is shown in Fig. 5(a). Since the vorticity of the vector potential ×$\nabla\times\mathbf{A}$ corresponds to the magnetic field, one can calculate the local flux in the Haldane model by counting the number of winding of the arrows in the figure. Then, the flux with a strength $-B$ appears around each lattice point, while each hexagonal plaquette has $2B$ flux inside. Therefore, the sum of the magnetic fluxes in each unit cell is zero, which means the absence of the external magnetic field. The internal magnetic field, however, breaks the time-reversal symmetry as one can see by confirming that the reversal of the arrows does not reproduce the same figure as in Fig. 5(a), which leads to the chiral edge current of the quantum anomalous Hall effect.
@@ 3.2.2 Band topology protected by symmetries
Figure 6: The band structure and phase diagram of the Kane-Mele model. The topological phase (a) exhibits gapless dispersions. The colors of gapless bands show at which side the gapless modes are localized. In contrast, the nontrivial phase (b) exhibits a gapped band structure, and thus topological modes are absent. These figures are adapted from C. L. Kane and E. J. Mele, “$Z_{2}$ Topological Order and the Quantum Spin Hall Effect” Phys. Rev. Lett. 95, 146802 (2005) [20], Copyright ©  2005 by the American Physical Society.
While the quantum Hall effect requires the breakdown of the time-reversal symmetry, the presence of symmetries can enrich the topological phases of matter. One typical example of such symmetry-protected topological phenomena is the quantum spin Hall effect [19, 20]. In the quantum spin Hall systems, one assumes the time-reversal symmetry, which implies that the exchange between spin-up and spin-down electrons only reverses the direction of motion (this situation is natural in materials without internal and external magnetic fields). Then, both spin-up and spin-down electrons can be localized at the edge of the sample and exhibit helical edge currents that flow in the opposite direction. Without symmetry protections, perturbative interactions between different spins can open the energy gap and break the gapless modes, while the time-reversal symmetry protects edge modes because it guarantees that edge current and its time-reversal counterpart must appear in a pair.
@@
where $C$ introduces interactions between spin-up and spin-down electrons that preserve the time-reversal symmetry. Calculating the dispersion relation of this Hamiltonian under the open boundary condition in the $x$ or $y$ direction, one can confirm the existence of a pair of gapless edge modes. Since each gapless band supports a unidirectional current along the edge of the sample, one can observe a helical current under the open boundary conditions, which is a characteristic behavior of the quantum spin Hall effect. We note that the spins of electrons may not be fully polarized at each gapless band because of the spin coupling term $C$, and thus the spin Hall conductivity is not quantized. However, both the quantum Hall effect and quantum spin Hall effect still share similar topological origins and topological edge modes.
@@
We shall explain the mechanism of the symmetry protection in the quantum spin Hall effect in more detail. Using a unitary matrix $T$, the time-reversal symmetry is described as $TH(\mathbf{k})T^{-1}=H^{\ast}(-\mathbf{k})$ with $H(\mathbf{k})$ being the Bloch Hamiltonian. If we focus on time-reversal invariant momenta where $\mathbf{k}_{\rm TRIM}$ and $-\mathbf{k}_{\rm TRIM}$ are equivalent in the Brillouin zone, the eigenvalues of $H(\mathbf{k}_{\rm TRIM})$ must be degenerate, which is known as the Kramers degeneracy. Thus, dispersions of a pair of gapless edge modes must cross at a time-reversal invariant momentum as shown in Fig. 7, which protects the edge modes from gap opening under continuous deformations of band structures. Meanwhile, when the band structures have two pairs of gapless modes, they can open a bandgap at wavenumbers which are variant under the time reversal. Therefore, topology of the quantum spin Hall system depends on whether the number of gapless modes is odd or even [144]. This $Z_{2}$-type classification is different from the topological classification of quantum Hall systems that are simply characterized by the number of gapless modes.
Figure 8: Edge modes and corner modes in the SSH model and the BBH model. (a) When we consider the fully dimerized case ($t_{1}=0$) in the SSH model, we can analytically obtain a zero mode localized at the left edge site surrounded by the red square. (b) In the BBH model, a corner mode appears at the site surrounded by the red square.
Topological insulators can be found in one-dimensional systems under another type of symmetry called chiral symmetry (sometimes called sublattice symmetry). The Su-Schriefer-Heeger (SSH) model is a prototypical model of such one-dimensional topological insulators, which was originally introduced to explain the electronic properties of polyacetylene [145]. The model is a lattice model whose Hamiltonian is described as
@@
where $\rm H.c.$ represents the Hermitian conjugate. In general, the chiral symmetry is defined as $\Gamma H({k})\Gamma^{-1}=-H^{\dagger}({k})$ with $H({k})$ being the Bloch Hamiltonian, and $H(k)$ and $\Gamma$ are described under a proper unitary transformation as
$H\left(k\right)=\left(\begin{array}[]{cc}0&h^{\dagger}(k)\\ h(k)&0\end{array}\right),$ (38)
$\Gamma=\left(\begin{array}[]{cc}I&0\\ 0&-I\end{array}\right),$ (41)
where $I$ represents an identity matrix. In the SSH model, $h(k)$ becomes a scalar function $h(k)=t_{1}+t_{2}e^{ik}$. Then, the topological invariant is the winding number of $h(k)$,
[…]
@@
[…]
@@
In the last part of this section, we shall briefly discuss the classification of the symmetry-protected topological phases. To discuss such topological classification, we need to introduce the particle-hole symmetry, which is defined as $CH(\mathbf{k})C^{-1}=-H^{T}(-\mathbf{k})$ where $H(\mathbf{k})$ is a Bloch Hamiltonian and $C$ is a unitary charge-conjugate operator. Based on the particle-hole, time-reversal, and chiral symmetry, Altland-Zirnbauer symmetry classes [148] classify the Hermitian system into 10 classes. Then, the classification table of topological phases [149, 150] indicates the existence or absence of topological phases in each symmetry class and dimension of the system. The classification table also tells us which types of topological invariants classify the topological phases; examples include the integer type such as the Chern number and the $Z_{2}$ type found in the quantum spin Hall systems. The winding number discussed in the previous paragraph is also included in the classification table as the $Z$ classification of the AIII class in one dimension. If we further consider spatial symmetries such as rotational symmetries, topological phases can be enriched [151, 152]. Another advantage of introducing spatial symmetries is that one can calculate symmetry indicators [144, 153], whose calculation is simpler than conventional topological invariants.

### v2

@@ 3.1 Band theory and its classical counterpart
Assuming periodically oscillating state $\psi(\mathbf{r};t)=e^{-iEt}\psi(\mathbf{r})$, one also obtains the time-independent Schrödinger equation $E\psi(\mathbf{r})=H(\mathbf{r})\psi(\mathbf{r})$, which is an eigenvalue problem of the linear operator $H$. In the following, we focus on this time-independent Schrödinger equation and discuss topological structures of eigenvalues and eigenvectors.
@@
${:=}$ $\left[-\frac{\hbar^{2}}{2m}(\nabla+i\mathbf{k})^{2}+V(\mathbf{r})\right]u_{\mathbf{k}}(\mathbf{r}),$
@@
Techniques of the Bloch Hamiltonian and the band structures are also useful in classical physics including active fluid. Basically, the Bloch Hamiltonian corresponds to the matrix obtained via the linear stability analysis of the Fourier modes in fluids [133]. Then, its dispersion relation can be considered as a band structure of fluid. In particular, when all the Fourier modes are marginally stable, one obtains real dispersion relation and thus can assume the matrix obtained via the linear stability analysis as an effective (quasi-)Hermitian Hamiltonian. Recent studies [22, 23, 134, 135, 136] have utilized the notions in condensed matter physics to develop metamaterials that control the dynamics of photonic and phononic systems. Metamaterials typically possess a periodic structure of waveguides, which has a clear analogy to the periodic alignment of atoms in crystalline solids. Utilizing this method to connect the linear stability analysis of fluid and condensed matter physics, topological active matter is also proposed as we discuss later in Sec. 5.
[…]
We also note that there is a subtle difference between quantum and classical systems when one considers the statistical properties of quantum many-body systems. In fermionic systems such as electrons in solids, each eigenstate is occupied by at most one electron because of Pauli’s exclusion principle. Then, in a ground state of the many-body system, energy bands are occupied from the lowest-energy one, and the highest energy occupied is called the Fermi energy. In conventional topological insulators, we consider the case that the band structure opens a bulk band gap at the Fermi energy and discuss the existence of gapless edge modes at that gap as we discuss in the following. In contrast, classical systems cannot define the Fermi energy due to the absence of the fermionic statistics. Instead, one can discuss the existence of topological gapless modes in any bulk band gaps. To observe such topological gapless modes, one should excite them by applying oscillating external force with corresponding frequencies.
@@ 3.2.1 Quantum Hall effect
Topology is a property of matter that is invariant under a continuous deformation. For example, considering a genus, the number of holes, we cannot either increase or decrease the genus without cutting the geometry or sticking different surfaces together. Thus, the genus is a topological invariant, and we cannot continuously transform objects with different genera into each other, as exemplified by a ball and a donut. Condensed matter physics employs the notion of topology into a wavenumber space by considering the topology of band structures [137, 138]. To discuss the topology of band structures, one must define the prohibited deformation of the Hamiltonian. In topological insulators, we first consider an energy gap at the Fermi energy, where no bands exist. Then, we exclude the deformation under which band structures cross the Fermi energy. Under such continuous deformation (i.e., adiabatic deformation), the bulk is always insulating. Then, one can find defective structures in a wavenumber space, and the number of such defects must be preserved under the continuous deformation in a similar manner as in the preservation of the genus. Thus, insulating phases of matter can be classified by such a topological number, which leads to the notion of topological insulators. Below we shall delineate the definition and physical meanings of the defective structures in band structures.
Figure 4: Quantum Hall effect. (a) Schematic of the quantum Hall effect. Cycrotolon motions of electrons are prevented at the edge of the sample, and instead, the skipping orbit can emerge, which induces the chiral edge current. (b) Experimental results of the quantized Hall voltage. The inset shows the top view of the device where the quantized Hall conductance was first observed. Panel (b) is adapted from K. v. Klitzing, G. Dorda, and M. Pepper, “New Method for High-Accuracy Determination of the Fine-Structure Constant Based on Quantized Hall Resistance” Phys. Rev. Lett. 45, 494 (1980) [15], Copyright © 1980 by the American Physical Society.
A prototypical example of topological materials is a quantum Hall system, which was experimentally discovered by von Klitzing et al. in 1980 [15]. They measured the Hall conductivity in semiconductors under the existence of the external magnetic field, and found its quantization. We first introduce some basic notions to understand this experiment. Under a uniform magnetic field perpendicular to a (effectively) two-dimensional semiconductor $\mathbf{B}=(0,0,B)$, the Hamiltonian is described as
@@
${:=}$ $\left[-\frac{\hbar^{2}}{2m}\left(\nabla-\frac{ie}{\hbar c}\mathbf{A}(\mathbf{r})\right)^{2}+V(\mathbf{r})\right]\psi(\mathbf{r}).$
by using the vector potential $\mathbf{A}(\mathbf{r})=(0,Bx,0)$ (we here adopt the Landau gauge). Solving the eigenequation of this Hamiltonian, one obtains discrete eigenenergies like the Hamiltonian under a harmonic potential, which is called the Landau level. Meanwhile, a uniform magnetic field induces the Lorentz force perpendicular to the velocity of an electron. As one can infer from classical electromagnetics, electronic and Lorentz forces are balanced when the electron moves perpendicular to both of them, and thus the corresponding current, i.e., the Hall current occurs. The Hall conductivity is defined as the Hall current divided by the electric field $\sigma_{xy}=j_{y}/E_{x}$. As we will discuss below, since each Landau level contributes to the Hall conductivity in a quantized way, the quantum Hall effect appears.
Von Klitzing et al. tuned the Fermi energy by changing the gate voltage (voltage between the gate and source in the inset of Fig. 4(b)) and obtained the contributions of different Landau levels. If the Fermi energy exists between Landau levels, the longitudinal conductivity becomes zero, and the Hall conductivity takes a constant value until the Fermi energy is increased to the higher Landau level. Thus, the Hall conductivity exhibits plateaus as shown in Fig. 4. Furthermore, the Hall conductivity is proportional to the number of occupied Landau levels and thus is quantized.
Thouless et al. [16] revealed that the quantization of the Hall conductance originates from the nontrivial topology of the wave functions. Their theoretical analysis starts from the Hamiltonian of an electron under the external magnetic field in Eq. (27). Here, we assume that the Fermi energy exists in the band gap. By considering the uniform electronic field $\vec{E}=(E,0)$ as a perturbation and using the Kubo formula, one can calculate the Hall current $\langle j_{y}\rangle=\langle e[\hat{y},H]\rangle$ at zero temperature [139]. This leads to the following expression of the Hall conductance
@@
While the Chern number and the quantized Hall conductance are the bulk properties of topological materials, nontrivial band topology also accompanies unconventional localized modes at the edge of the sample. The appearance of such edge modes can be intuitively understood as follows; since electrons under a uniform magnetic field exhibit cyclotron motions, each electron circularly moves in the bulk of the quantum Hall system. Meanwhile, near the boundary of the sample, the circular motions are interfered by the potential barrier, and instead, electrons bounce at the boundary. The iteration of the curved motions and the bounds leads to a skipping orbit of an electron (cf. the blue curved allows in Fig. 4(a)), where electrons unidirectionally flow along the edge of the sample. Therefore, the quantum Hall system exhibits edge-localized modes that exhibit unidirectional edge flow. We note that a similar boundary flow can occur in rotating active particles, which is an origin of nontrivial topology in chiral active matter in Sec. 5.2.1.
@@
In general, the correspondence between the bulk topological invariant and the gapless boundary modes is believed as the bulk-boundary correspondence. To mathematically formulate this in two-dimensional systems, one considers the left semi-infinite boundary condition, under which the open boundary exists at $x=0$, and the system extends to the $x>0$ region and is periodic in the $y$ direction. Then, the bulk-boundary correspondence indicates
$C=n_{L}-n_{R},$ (33)
where $C$ is the sum of the Chern numbers of occupied bands in Eq. (29) and $n_{L}$ and $n_{R}$ are the numbers of gapless edge modes with left and right chirality (i.e., propagating in the upper and lower direction), respectively. One can also extend this bulk-boundary correspondence to general dimensions and various boundary conditions, and in many cases, the bulk-boundary correspondence is mathematically proved.
@@
where $\langle\mathbf{r},\mathbf{r}^{\prime}\rangle$ and $\langle\langle\mathbf{r},\mathbf{r}^{\prime\prime}\rangle\rangle$ represent the pairs of nearest and next-nearest neighbor sites, respectively, and $\rm H.c.$} represents the Hermitian conjugate of the previous term. $t_{1}$, $t_{2}$ are real-valued parameters determining the hopping amplitudes, and $\phi$ is the phase gained via the hopping to the next-nearest site. $\sigma(\mathbf{r})$ determines the sign of the on-site potential and depends on the sublattice. Such phase factor plays the role of a vector potential (in other words, a gauge field) in a lattice system, and the direction of the effective vector potential is shown in Fig. 5(a). Since the vorticity of the vector potential $\nabla\times\mathbf{A}$ corresponds to the magnetic field, one can calculate the local flux in the Haldane model by counting the number of winding of the arrows in the figure. Then, the flux with a strength B-B appears around each lattice point, while each hexagonal plaquette has $2B$ flux inside. Therefore, the sum of the magnetic fluxes in each unit cell is zero, which means the absence of the external magnetic field. The internal magnetic field, however, breaks the time-reversal symmetry as one can see by confirming that the reversal of the arrows does not reproduce the same figure as in Fig. 5(a), which leads to the chiral edge current of the quantum anomalous Hall effect.
@@
In this section, we have introduced key concepts in band topology, i.e., topological invariants such as the Chern number and their bulk-boundary correspondence. We note that both the Chern number and corresponding edge modes are properties of eigenvectors and thus are not related to the quantumness of condensed matter. Thus, one can extend these notions to classical systems including active matter by considering effective Hamiltonians via the linear stability analysis. We will discuss this point later in Sec. 5.
@@ 3.2.2 Band topology protected by symmetries
Figure 6: The band structure and phase diagram of the Kane-Mele model. The inset shows the phase diagram of the model, where the vertical axis represents the nearest neighbor Rashba term that introduces spin couplings and breaks mirror symmetry, and the horizontal axis represents a staggered potential. The topological phase (a) exhibits gapless dispersions. The colors of gapless bands show at which side the gapless modes are localized. In contrast, the nontrivial phase (b) exhibits a gapped band structure, and thus topological modes are absent. These figures are adapted from C. L. Kane and E. J. Mele, $Z_{2}$ Topological Order and the Quantum Spin Hall Effect” Phys. Rev. Lett. 95, 146802 (2005) [18], Copyright © 2005 by the American Physical Society.
While the quantum Hall effect requires the breakdown of the time-reversal symmetry by magnetic flux because of the chiral edge current, the presence of symmetries can enrich the topological phases of matter. One typical example of such symmetry-protected topological phenomena is the quantum spin Hall effect [17, 18]. In the quantum spin Hall systems, one assumes the time-reversal symmetry, which implies that the exchange between spin-up and spin-down electrons only reverses the direction of motion (this situation is natural in materials without internal and external magnetic fields). Then, both spin-up and spin-down electrons can be localized at the edge of the sample and exhibit helical edge currents that flow in the opposite direction. Without symmetry protections, perturbative interactions between different spins can open the energy gap and break the gapless modes, while the time-reversal symmetry protects edge modes because it guarantees that edge current and its time-reversal counterpart must appear in a pair.
@@
where $C$ introduces interactions between spin-up and spin-down electrons that preserve the time-reversal symmetry. Calculating the dispersion relation of this Hamiltonian under the open boundary condition in the $x$ or $y$ direction, one can confirm the existence of a pair of gapless edge modes localized at the same side of the open system (cf. Fig. 6). Since each gapless band supports a unidirectional current along the edge of the sample, one can observe a helical current under the open boundary conditions, which is a characteristic behavior of the quantum spin Hall effect. We note that the spins of electrons may not be fully polarized at each gapless band (i.e., the eigenstate at each gapless band is not an eigenstate of the spin operator $S_{z}$) because of the spin coupling term $C$, and thus the spin Hall conductivity is not quantized. However, both the quantum Hall effect and quantum spin Hall effect still share similar topological origins and topological edge modes.
@@
We shall explain the mechanism of the symmetry protection in the quantum spin Hall effect in more detail. Using a unitary matrix $T$, the time-reversal symmetry is described as $TH(\mathbf{k})T^{-1}=H^{\ast}(-\mathbf{k})$ with $H(\mathbf{k})$ being the Bloch Hamiltonian. If we focus on time-reversal invariant momenta where $\mathbf{k}_{\rm TRIM}$ and $\mathbf{k}_{\rm TRIM}$ are equivalent in the Brillouin zone, the eigenvalues of $H(\mathbf{k}_{\rm TRIM})$ must be degenerate, which is known as the Kramers degeneracy. Thus, dispersions of a pair of gapless edge modes must cross at a time-reversal invariant momentum as shown in Fig. 7, which protects the edge modes from gap opening under continuous deformations of band structures. Meanwhile, when the band structures have two pairs of gapless modes, they can open a bandgap at wavenumbers which are variant under the time reversal. Therefore, topology of the quantum spin Hall system depends on whether the number of gapless modes is odd or even [149]. This $Z_{2}$-type classification is different from the topological classification of quantum Hall systems that are simply characterized by the number of gapless modes.
Figure 8: Edge modes and corner modes in the SSH model and the BBH model. The black lines and dotted lines represent the nearest neighbor hoppings with strengths $t_{2}$ and $t_{1}$, respectively. (a) When we consider the fully dimerized case ($t_{1}=0$) in the SSH model, we can analytically obtain a zero mode localized at the left edge site surrounded by the red square. (b) In the BBH model, a corner mode appears at the site surrounded by the red square.
Topological insulators can be found in one-dimensional systems under another type of symmetry called chiral symmetry (sometimes called sublattice symmetry). The Su-Schriefer-Heeger (SSH) model is a prototypical model of such one-dimensional topological insulators, which was originally introduced to explain the electronic properties of polyacetylene [150] and later used as a platform to study fundamental theories of topological insulators (see Sec. 4.3.3 examples of studies in the non-Hermitian topology). The model is a lattice model whose Hamiltonian is described as
@@
where $\rm H.c.$ represents the Hermitian conjugate. In general, the chiral symmetry is defined as $\Gamma H({k})\Gamma^{-1}=-H^{\dagger}({k})$ with H(k)H({k}) being the Bloch Hamiltonian, and $H(k)$ and $\Gamma$ are described under a proper unitary transformation as
$H\left(k\right)=\left(\begin{array}[]{cc}O&h^{\dagger}(k)\\ h(k)&O\end{array}\right),$ (39)
$\Gamma=\left(\begin{array}[]{cc}I&O\\ O&-I\end{array}\right),$ (42)
where $I$ represents an identity matrix, and $O$ represents a zero matrix. In the SSH model, $h(k)$ becomes a scalar function $h(k)=t_{1}+t_{2}e^{ik}$. Then, the topological invariant is the winding number of $h(k)$,
$w=\frac{1}{2\pi i}\int_{0}^{2\pi}dk\frac{d}{dk}\log(\det h(k)).$ (43)
@@
$\frac{\phi}{2\pi}=\frac{1}{2\pi i}\int_{0}^{2\pi}dk\,\vec{u}(k)\cdot\left[\partial_{k}\vec{u}(k)\right]$ (44)
@@
In the last part of this section, we shall briefly discuss the classification of the symmetry-protected topological phases. To discuss such topological classification, we need to introduce the particle-hole symmetry, which is defined as $CH(\mathbf{k})C^{-1}=-H^{T}(-\mathbf{k})$ where $H(\mathbf{k})$ is a Bloch Hamiltonian and $C$ is a unitary charge-conjugate operator. Based on the particle-hole, time-reversal, and chiral symmetry, Altland-Zirnbauer (AZ) symmetry classes [153] classify the Hermitian system into 10 classes. Then, the classification table of topological phases [154, 155] indicates the existence or absence of topological phases in each symmetry class and dimension of the system. The classification table also tells us which types of topological invariants classify the topological phases; examples include the integer type such as the Chern number and the $Z_{2}$ type found in the quantum spin Hall systems. The winding number discussed in the previous paragraph is also included in the classification table as the $Z$ classification of the AIII class in one dimension. If we further consider spatial symmetries such as rotational symmetries, topological phases can be enriched [156, 157]. Another advantage of introducing spatial symmetries is that one can calculate symmetry indicators [149, 158], whose calculation is simpler than conventional topological invariants.
We note that most of classical linear dynamics are described as $\partial_{t}\psi=A\psi$ by using a real matrix $A$ and real state vectors $\psi$. By assuming $iA$ as the effective Hamiltonian $H$, such a linear dynamics must have the particle-hole symmetry $H=-H^{\ast}$. The time-reversal and chiral symmetries also widely exist by using only reciprocal elements or preparing time-reversal counterparts (effective spins). Therefore, symmetry-protected topological phases are ubiquitous in classical systems such as active matter. In fact, we will discuss the non-Hermitian topology of active matter protected by the chiral symmetry in Sec. 5.3.